# Knockoffs for low dimensions: changing the nominal level post-hoc to gain power while controlling the FDR

### v1

@@ Abstract
Knockoffs are a powerful tool for controlled variable selection with false discovery rate (FDR) control. However, while they are frequently used in high-dimensional regressions, they lack power in low-dimensional and sparse signal settings. One of the main reasons is that knockoffs require a minimum number of selections, depending on the nominal FDR level. In this paper, we leverage e-values to allow the nominal level to be switched after looking at the data and applying the knockoff procedure. In this way, we can increase the nominal level in cases where the original knockoff procedure does not make any selections to potentially make discoveries. Also, in cases where the original knockoff procedure makes discoveries, we can often decrease the nominal level to increase the precision. These improvements come without any costs, meaning the results of our post-hoc knockoff procedure are always more informative than the results of the original knockoff procedure. Furthermore, we apply our technique to recently proposed derandomized knockoff procedures.
@@ 1 Introduction
In exploratory analyses, the identification of relevant variables is a critical step toward understanding complex data-generating mechanisms and uncovering meaningful associations. [^1] We refer to the task as variable selection, noting that equivalent terminology includes feature selection, biomarker discovery, or covariate identification, depending on the application domain. Over the past years, a wide range of methods that leverage modern machine learning techniques have been proposed to quantify the importance of candidate variables. While these approaches provide powerful tools for ranking variables according to their apparent influence, they typically lack rigorous guarantees for controlled variable selection. In particular, translating variable importance scores into a selected subset of variables while controlling false discoveries remains a major methodological challenge. To bridge this gap, recent methods have aimed to develop inference procedures that combine statistical rigor with modeling flexibility, with the knockoff framework emerging as a particularly effective approach [1, 4].
@@ 1.1 Controlled variable selection and knockoffs
$H_{0}^{i}:X_{i}\perp\!\!\!\perp Y\mid\mathbf{X}_{-i},$ (2)
@@
$I_{0}=\{i:H_{0}^{i}\text{ is true}\},\qquad I_{1}=[p]\setminus I_{0}.$ (3)
A variable selection procedure outputs a set of discoveries/rejections $R\subseteq[p]$, and its quality is measured by the false discovery rate
$\mathrm{FDR}=\mathbb{E}\!\left[\frac{|R\cap I_{0}|}{|R|\vee 1}\right],$ (4)
@@
The task of identifying the relevant variables, that is, those in $I_{1}$, naturally leads to the question of how to design a variable selection procedure that can recover as many non-nulls as possible while rigorously controlling false discoveries. The knockoff framework of Barber and Candès [1], refined and generalized by Candès et al. [4], addresses this goal by enabling controlled variable selection: it yields a data-driven rejection set of putative discoveries while provably controlling the false discovery rate (FDR) at a user-specified level $\alpha\in(0,1)$, meaning $\textnormal{FDR}\leq\alpha$. Methods based on knockoffs construct synthetic variables that act as negative controls, enabling data-adaptive selection with provable FDR guarantees, while making minimal assumptions on the conditional distribution of $Y\mid\mathbf{X}$. A full description of the knockoff construction and filtering procedure is deferred to Section 2.1.
@@ 1.2 Motivation of our work
While the knockoff framework was primarily developed for high-dimensional problems, it is equally valuable to have methods that provide controlled variable selection in low- and moderate-dimensional settings, particularly when the underlying signal is expected to be sparse. In many scientific applications, the number of truly relevant variables is small, even if a large number of variables are measured. When identifying effect modifiers (i.e., predictive variables) in clinical trials, the total number of available clinical variables is typically modest—often on the order of 10–50 patient characteristics—reflecting the scale of routinely collected baseline data. Within this set, it is generally assumed that only a small subset truly drives heterogeneity in treatment effect, leading to a sparse predictive structure in most methodological developments and simulation studies [16, 24, 17, 30]. Interestingly, in such settings, the vanilla knockoff procedure exhibits certain limitations that motivate further methodological development. In our work we will focus on two main limitations of the original knockoffs:
The minimum number of discoveries required to achieve statistical significance is determined by the pre-specified FDR level $\alpha$, and it is $1/\alpha,$ which can be overly conservative when only a few signals are present. As a result, the procedure may fail to flag any discoveries even when a small but genuine signal exists.
@@ 1.3 Contribution
E-values [19, 20] are a novel tool for hypothesis testing that has proven particularly useful for sequential anytime-valid inference [7, 27, 34, 9], testing of composite nulls in irregular models [33, 15], multiple testing [32, 11, 31, 36], and, most recently, for hypothesis testing at data-dependent nominal levels [8, 14, 6, 36]. In this paper, we exploit the two latter aspects of e-values to construct a knockoff procedure that allows to switch the nominal FDR level post-hoc (see Section 2.2 for a detailed description of FDR control with data-dependent $\alpha$). This improves the original knockoff method [1, 4] in two ways.
@@
Figure 1 illustrates the benefits of the proposed post-hoc adjustment. In Figure 1(a), we observe that the post-hoc procedure dramatically increases power: across all settings for the number of relevant variables, our method successfully identifies nearly all relevant variables (i.e., power $\approx 1$). In contrast, the original knockoff procedure exhibits very low power when the number of relevant variables is small. This is primarily because the minimum number of discoveries is determined by the nominal FDR level $\alpha$, which in our case corresponds to $1/\alpha=5$. To illustrate this, when the actual number of relevant variables is 3, the original procedure produces no discoveries in 1542 out of 2000 iterations and exactly 5 discoveries in 229 iterations, which are the two most frequent outcomes. By contrast, the post-hoc method most frequently selects 3 variables (1002 iterations) and 4 variables (467 iterations), corresponding to the two most common outcomes, closely matching the true number of relevant variables.
The gains in power come with only a small sacrifice in the (averaged) nominal $\alpha$ value, as illustrated in Figure 1(b) for scenarios with a low number of relevant variables (3 and 4). Moreover, our post-hoc adjustment can reduce the effective nominal level $\alpha$ when the problem is easier than anticipated, while maintaining the same power as the original procedure. It should be noted that our improvement is a “free-lunch”, meaning our approach never makes less discoveries and it only reports a larger significance level $\alpha$ in cases where the original method did not make any rejections. Therefore, the results reported by our method are always more informative.
@@
Figure 1: This figure illustrates the main contribution of our work, namely that the proposed post-hoc adjustment of the $\alpha$ values substantially increases statistical power compared to the original approach. For small values of $p_{\text{relevant}}$, this improvement comes at a minor cost of slightly inflated $\alpha$ values, whereas for larger $p_{\text{relevant}}$, the post-hoc method can even reduce $\alpha$ while maintaining comparable power. The nominal level of $\alpha$ for the original KO method is set to $0.20.$ A full description of the data-generating process, parameter choices, and simulation protocol is provided in Section 5.
@@
Since the knockoff procedure is based on the random generation of knockoff variables, the results are heavily exposed to external randomness. In order reduce its dependence on the random data generator, Ren et al. [22], Ren and Barber [21] introduced derandomized knockoff procedures by sampling multiple knockoff copies for each variable. These derandomized knockoff procedures require the pre-specification of an additional parameter, on which the performance and guarantee of the procedure depend considerably. Ren et al. [22], Ren and Barber [21] proposed heuristics, simulations and data-splitting methods to obtain a reasonable parameter value. We show that the parameter can be chosen arbitrarily based on the entire data, thus increasing the efficiency and utility of the derandomized methods enormously. Furthermore, we show how the derandomized procedures can be further uniformly improved.
@@ 1.4 Outline
In the following, we start to briefly summarize the knockoff procedure by Barber and Candès [1], Candès et al. [4] and the e-Closure Principle to derive FDR procedures with post-hoc error control by Xu et al. [36]. Afterwards, in Section 3, we introduce our post-hoc knockoff procedure and derive its error control. In Section 4, we consider the derandomized procedures and our proposed improvements. Finally, in Section 5, we provide extensive simulation results that illustrate the superiority of our methods. [^2] The code implementing the methods described in this paper is publicly available at https://github.com/sechidis/posthoc_knockoffs.
@@ 2.1 The BC-knockoff procedure
The knockoff procedure introduced by Barber and Candès [1] and extended by Candès et al. [4] is a powerful framework for variable selection while controlling the false discovery rate (FDR). Since we consider multiple variations of their initial approach in this paper, we call it BC-knockoff procedure in the following. The strategy is composed of three steps:
@@
    $\tilde{\mathbf{X}}\perp\!\!\!\perp Y\,|\,\mathbf{X}.$ (5)
    This property is straightforwardly guaranteed as long as neither the response $Y$ in the knockoff generation process. Its purpose is to maintain neutrality during model fitting and selection steps.
    Various algorithms have been proposed to construct knockoffs that obey the exchangability property [26, 23, 13, 2, 29, 10].
@@
    where $Z_{i}$ and $\tilde{Z}_{i}$ capture importance for the original variable $X_{i}$ and its knockoff $\tilde{X}_{i}$, respectively. The function $q$ represents an appropriate model or statistical procedure that assigns importance scores. These statistics are constructed so that swapping any variable with its knockoff results in the corresponding components of the statistic being swapped as well. Most variable importance score estimation procedures, such as those based on regression coefficients and permutation importance measures, inherently satisfy this symmetry property [4]. This exchangeability is a key requirement for maintaining the validity of knockoffs.
@@
    $f(u,v)=u-v,\quad\text{or}\quad f(u,v)=|u|-|v|.$ (6)
@@
    The last years, many powerful knockoff statistics have been introduced in the literature, see for example Spector and Fithian [28]. Furthermore, there are works that derive knockoff statistics targeting effect modification in randomized trials: for example, Sechidis et al. [25] develop predictive knockoff statistics that quantify the extent to which baseline variables modify treatment effects in RCTs, enabling FDR-controlled selection of predictive variables (i.e., effect modifiers) rather than purely prognostic ones.
@@
    $R^{\textnormal{BC}}\coloneqq\{i\in[p]:W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}\},\quad\text{where }T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}\coloneqq\inf\left\{t>0:\frac{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq t\}}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}}\geq 1/\alpha^{\textnormal{kn}}\right\}.$ (7)
@@
With the latter result, one can show that the BC-knockoff procedure controls the FDR at level $\alpha=\alpha^{\textnormal{kn}}$ [1, 4]. Similar to [21], we distinguish between the FDR level $\alpha$ and the parameter $\alpha^{\textnormal{kn}}$ used for the knockoff stopping time in (7), since they can deviate later in the paper. Also, in this paper we propose to choose the significance level $\alpha$ post-hoc, however, the parameter $\alpha^{\textnormal{kn}}$ needs to be specified in advance and can be interpreted as an initial or targeted significance level.
@@ 2.2 Post-hoc choice of $\alpha$
The Closure Principle of Marcus et al. [18] has long been known as a fundamental approach to multiple testing with familywise error rate (FWER) control. Recently, Xu et al. [36] introduced an e-Closure Principle for false discovery rate (FDR) control. Given a local e-value $E_{S}$ for each intersection hypothesis $H_{0}^{S}=\bigcap_{i\in S}H_{0}^{i}$, $S\subseteq[p]$, meaning $\mathbb{E}_{H_{0}^{S}}[E_{S}]\leq 1$, we consider the closure set
@@
where $\boldsymbol{E}=(E_{S})_{S\subseteq[p]}$ and $\textnormal{FDP}_{S}(R)=\frac{|S\cap R|}{|R|\lor 1}$. Xu et al. [36] showed that any rejection set $R\in\mathcal{R}_{\alpha}(\boldsymbol{E})$ controls the FDR and any procedure that controls the FDR at level $\alpha$ can be constructed in this way.
Furthermore, Xu et al. [36] showed that these closure sets even permit a more sophisticated error rate control that goes beyond the classical notion of FDR control — $\alpha$ can be chosen post-hoc! Given a family of e-values $\boldsymbol{E}=(E_{S})_{S\subseteq[p]}$, they proved that
$\mathbb{E}\left[\sup_{\alpha\in(0,1]}\max_{R\in\mathcal{R}_{\alpha}(\boldsymbol{E})}\frac{\textnormal{FDP}_{I_{0}}(R)}{\alpha}\right]\leq 1.$ (9)
@@
Note that the post-hoc error rate control in (10) is slightly different from classical FDR control (where the FDR is bounded by $\alpha$), since $\tilde{\alpha}$ is inside the expectation. If $\tilde{\alpha}$ was fixed in advance, then the two notions would be equivalent due to the linearity of the expected value. However, since $\tilde{\alpha}$ is allowed to be a random variable depending on the data, the classical notion needs to be adapted by drawing $\tilde{\alpha}$ inside the expectation. This notion of post-hoc validity was also considered in [8, 14] for single hypotheses.
@@
Example 1.
Suppose all null hypotheses are true ($I_{0}=[p]$). In that case $\textnormal{FDP}_{I_{0}}(R)$ becomes $1$ if $R\neq\emptyset$ and 0 otherwise. Suppose we have a multiple testing procedure $R_{\alpha}$ for each $\alpha\in(0,1]$ that has exact error rate control, meaning $\mathbb{E}[\textnormal{FDP}_{I_{0}}(R_{\alpha})]=\alpha$ (for example, this is satisfied by the Benjamini-Hochberg procedure [3] with independent and uniformly distributed null p-values). Now set $\tilde{\alpha}=\inf\{\alpha\in(0,1]:R_{\alpha}\neq\emptyset]$. Then, we have

### v2

# Choosing the nominal level post-hoc with knockoffs using e-values

@@ Abstract
The knockoff filter is a powerful tool for controlled variable selection with false discovery rate (FDR) control. In this paper, we leverage e-values to allow the nominal FDR level to be switched post-hoc, after looking at the data and applying the knockoff procedure. This approach addresses a significant limitation of standard knockoffs: while frequently used in high-dimensional regressions, they often lack power in low-dimensional and sparse signal settings. One of the main reasons for this is that the knockoff filter requires a minimum number of selections that depends strictly on the nominal FDR level. By utilizing e-values, we can increase the nominal level in cases where the original procedure makes no discoveries, or decrease it to improve precision when discoveries are abundant. These improvements come without any costs, meaning the results of our post-hoc procedure are always more informative than those of the original knockoff filter. We extend this methodology to recently proposed derandomized knockoff procedures and demonstrate its utility in variable selection problems relevant to drug development using real clinical trial data.
Keywords: controlled variable selection, knockoffs, e-values
@@ 1 Introduction
In exploratory analyses, the identification of relevant variables is a critical step toward understanding complex data-generating mechanisms and uncovering meaningful associations. [^1] We refer to the task as variable selection, noting that equivalent terminology includes feature selection, biomarker discovery, or covariate identification, depending on the application domain. Over the past years, a wide range of methods that leverage modern machine learning techniques have been proposed to quantify the importance of candidate variables. While these approaches provide powerful tools for ranking variables according to their apparent influence, they typically lack rigorous guarantees for controlled variable selection. In particular, translating variable importance scores into a selected subset of variables while controlling false discoveries remains a major methodological challenge. To bridge this gap, recent methods have aimed to develop inference procedures that combine statistical rigor with modeling flexibility, with the knockoff framework emerging as a particularly effective approach (Barber and Candès, 2015; Candès et al., 2018).
@@ 1.1 Controlled variable selection and knockoffs
$H_{0}^{i}:X_{i}\perp\!\!\!\perp Y\mid\mathbf{X}_{-i},$
@@
$I_{0}=\{i:H_{0}^{i}\text{ is true}\},\qquad I_{1}=[p]\setminus I_{0}.$
A variable selection procedure outputs a set of discoveries/rejections $R\subseteq[p]$, and its quality is measured by the false discovery rate (FDR), which is defined as the expected number of false discovery proportion (FDP):
$\mathrm{FDR}=\mathbb{E}\!\left[\mathrm{FDP}\right],~~\mathrm{with}~~\mathrm{FDP}=\frac{|R\cap I_{0}|}{|R|\vee 1},$
@@
The task of identifying the relevant variables, that is, those in $I_{1}$, naturally leads to the question of how to design a variable selection procedure that can recover as many non-nulls as possible while rigorously controlling false discoveries. The knockoff framework of Barber and Candès (2015), refined and generalized by Candès et al. (2018), addresses this goal by enabling controlled variable selection: it yields a data-driven rejection set of putative discoveries while provably controlling the FDR at a user-specified level $\alpha\in(0,1)$, meaning $\textnormal{FDR}\leq\alpha$. Methods based on knockoffs (KO) construct synthetic variables that act as negative controls, enabling variable selection with provable FDR guarantees, while making minimal assumptions on the conditional distribution of $Y\mid\mathbf{X}$. A full description of the knockoff construction and filtering procedure is deferred to Section 2.1.
@@ 1.2 Motivation of our work
While the knockoff framework was primarily developed for high-dimensional problems, it is equally valuable to have methods that provide controlled variable selection in low- and moderate-dimensional settings, particularly when the underlying signal is expected to be sparse. In many scientific applications, the number of truly relevant variables is small, even if a large number of variables are measured. When identifying effect modifiers (i.e., predictive variables) in clinical trials, the total number of available clinical variables is typically modest, often on the order of 10–50 patient characteristics, reflecting the scale of routinely collected baseline data. Within this set, it is generally assumed that only a small subset truly drives heterogeneity in treatment effect, leading to a sparse predictive structure in most methodological developments and simulation studies (Lipkovich et al., 2017; Sechidis et al., 2018; Lipkovich et al., 2024; Sun et al., 2024). Interestingly, in such settings, the vanilla knockoff procedure exhibits certain limitations that motivate further methodological development. In our work we will focus on two main limitations of the original knockoffs:
The minimum number of discoveries required to achieve statistical significance is determined by the pre-specified FDR level $\alpha$, and it is $1/\alpha,$ which can be overly conservative when only a few signals are present (Zimmermann et al., 2024; Luo et al., 2025). As a result, the procedure may fail to flag any discoveries even when a small but genuine signal exists.
@@ 1.3 Contribution
E-values (Ramdas and Wang, 2024; Ramdas et al., 2023) are a novel tool for hypothesis testing that has proven particularly useful for sequential anytime-valid inference (Grünwald et al., 2024; Shafer, 2021; Waudby-Smith and Ramdas, 2020; Howard et al., 2021), testing of composite nulls in irregular models (Wasserman et al., 2020; Larsson et al., 2024), multiple testing (Wang and Ramdas, 2022; Ignatiadis et al., 2024; Vovk and Wang, 2021; Xu et al., 2025), and, most recently, for hypothesis testing at data-dependent nominal levels (Grünwald, 2024; Koning, 2023; Gauthier et al., 2025; Xu et al., 2025) (see Section 2.2 for a detailed description of FDR control with data-dependent $\alpha$). In this paper, we exploit the two latter aspects of e‑values to construct a knockoff procedure, described in Algorithm 1, that allows the nominal FDR level to be adjusted post‑hoc. This improves the original knockoff method (Barber and Candès, 2015; Candès et al., 2018) in two ways.
@@
Figure 1 illustrates the benefits of the proposed post‑hoc adjustment. In Figure 1a, we observe that the post-hoc procedure dramatically increases power: across all settings for the number of relevant variables, our method successfully identifies nearly all relevant variables (i.e., power $\approx 1$). In contrast, the original knockoff procedure exhibits very low power when the number of relevant variables is small. This is primarily because the minimum number of discoveries is determined by the nominal FDR level $\alpha$, which in our case corresponds to $1/\alpha=5$. To illustrate this, when the actual number of relevant variables is 3, the original procedure produces no discoveries in 1490 out of 2000 iterations and exactly 5 discoveries in 238 iterations, which are the two most frequent outcomes. By contrast, the post-hoc method most frequently selects 3 variables (924 iterations) and 4 variables (499 iterations), corresponding to the two most common outcomes, closely matching the true number of relevant variables.
The gains in power come with only a small sacrifice in the (averaged across simulations) nominal $\alpha$ value (solid lines), as illustrated in Figure 1b for scenarios with a low number of relevant variables (3 and 4). Interestingly, the average FDP acrosss the simulations (i.e., empirical FDR) remains below 0.20 for both methods (dashed lines). Moreover, our post-hoc adjustment can reduce the effective nominal level $\alpha$ when the problem is easier than anticipated, while maintaining the same power as the original procedure.
@@
Figure 1: This figure illustrates the main contribution of our work, namely that the proposed post-hoc adjustment of the $\alpha$ values substantially increases statistical power compared to the original approach. For small values of $p_{\text{relevant}}$, the $\alpha$ level is slightly increased to obtain rejections in cases where the original method does not make any discoveries, whereas for larger $p_{\text{relevant}}$, the post-hoc method can even reduce $\alpha$ while maintaining similar (or better) power. The nominal level of $\alpha$ for the original KO method is set to $0.20.$ Results are averaged over 2000 runs, and a full description of the data‑generating process, parameter choices, and simulation protocol is provided in Section 5.
It should be noted that our improvement is a “free-lunch”, meaning our approach never makes less discoveries and it only reports a larger significance level $\alpha$ in cases where the original method did not make any rejections. Therefore, the results reported by our method are always more informative. To further illustrate these properties, we examine the setting with 3 relevant variables, as detailed in Table 1. Out of the 1490 runs where the original method failed to yield any discoveries, the post-hoc approach successfully identified an average of 3.27 variables. While this was achieved at a higher average nominal level ($0.313$), it highlights the post-hoc method’s ability to extract information where the original method remains uninformative. Conversely, in the 510 runs where the original method successfully identified variables, our post-hoc approach matched the original discoveries, but did so while maintaining tighter error control, achieving an average nominal level of $0.178$ compared to the original $0.20$.
Table 1: Comparison of original vs posthoc KO number of discoveries and nominal $\alpha$ levels
Original KO at nominal level $\alpha=0.20$ predefined Post-hoc KO avg. number of discoveries Post-hoc KO avg. nominal level $\alpha$ post-hoc determined
No discoveries in 1490 runs out of 2000 3.27 0.313
$\geq 5$ discoveries in 510 runs out of 2000 Same as original KO 0.178
@@
Since the knockoff procedure is based on the random generation of knockoff variables, the results are heavily exposed to external randomness. In order to reduce dependence on the random data generator, Ren et al. (2023) and Ren and Barber (2024) introduced derandomized knockoff procedures by sampling multiple knockoff copies for each variable. Specifically, Ren et al. (2023) focuse on controlling the per-family error rate (PFER), and Ren and Barber (2024) provide a framework for FDR control. These derandomized procedures require the pre-specification of an additional parameter, on which the performance and guarantee of the procedure significantly depend. We show that the parameter can be chosen arbitrarily based on the entire data while maintaining valid error control, thus increasing the efficiency and utility of the derandomized methods enormously. Furthermore, we show how the derandomized procedures can be further uniformly improved.
@@ 1.4 Outline
In the following, we start to briefly summarize the knockoff procedure and the e-Closure Principle to derive FDR procedures with post-hoc error control. Afterwards, in Section 3, we introduce our post-hoc knockoff procedure and derive its error control. In Section 4, we consider the derandomized post-hoc procedures and our proposed improvements. In Section 5, we provide extensive simulation results that demonstrate the effectiveness of our approach compared to existing methods. [^2] The code implementing the methods described in this paper is publicly available at https://github.com/sechidis/posthoc_knockoffs. Finally, in Section 6, we illustrate how to use our proposed methods in clinical trial data, demonstrating its effectiveness in identifying prognostic and predictive variables, two tasks very critical in modern drug development.
@@ 2.1 The BC-knockoff procedure
The knockoff procedure introduced by Barber and Candès (2015) and extended by Candès et al. (2018) is a powerful framework for variable selection while controlling the FDR. Since we consider multiple variations of their initial approach in this paper, we call it BC-knockoff procedure in the following. The strategy is composed of three steps:
@@
    $\tilde{\mathbf{X}}\perp\!\!\!\perp Y\,|\,\mathbf{X}.$
    This property is automatically satisfied if the response $Y$ is not used in the knockoff generation process.
    Various algorithms have been proposed to construct knockoffs that obey the exchangability property (Sesia et al., 2018; Romano et al., 2019; Jordon et al., 2019; Bates et al., 2020; Huang and Janson, 2020; Spector and Janson, 2022).
@@
    where $Z_{i}$ and $\tilde{Z}_{i}$ capture importance for the original variable $X_{i}$ and its knockoff $\tilde{X}_{i}$, respectively. The function $q$ represents an appropriate model or statistical procedure that assigns importance scores. These statistics are constructed so that swapping any variable with its knockoff results in the corresponding components of the statistic being swapped as well. Most variable importance score estimation procedures, such as those based on regression coefficients and permutation importance measures, inherently satisfy this symmetry property (Candès et al., 2018). This exchangeability is a key requirement for maintaining the validity of knockoffs.
@@
    $f(u,v)=u-v,\quad\text{or}\quad f(u,v)=|u|-|v|.$
@@
    The last years, many powerful knockoff statistics have been introduced in the literature, see for example Spector and Fithian (2024). Furthermore, there are works that derive knockoff statistics targeting effect modification in randomized trials: for example, Sechidis et al. (2021) develop predictive knockoff statistics that quantify the extent to which baseline variables modify treatment effects in RCTs, enabling FDR-controlled selection of predictive variables (i.e., effect modifiers) rather than purely prognostic ones.
@@
    $R^{\textnormal{BC}}$ $\coloneqq\{i\in[p]:W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}\},$
    $\text{where }T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$ $\coloneqq\inf\left\{t>0:\frac{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq t\}}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}}\geq 1/\alpha^{\textnormal{kn}}\right\}.$ (1)
@@
With the latter result, one can show that the BC-knockoff procedure controls the FDR at level $\alpha=\alpha^{\textnormal{kn}}$ (Barber and Candès, 2015; Candès et al., 2018). Similar to Ren and Barber (2024), we distinguish between the FDR level $\alpha$ and the parameter $\alpha^{\textnormal{kn}}$ used for the knockoff stopping time in (1), since they can deviate later in the paper. Also, in this paper we propose to choose the significance level $\alpha$ post-hoc, however, the parameter $\alpha^{\textnormal{kn}}$ needs to be specified in advance and can be interpreted as an initial or targeted significance level.
@@ 2.2 Post-hoc choice of $\alpha$
The Closure Principle of Marcus et al. (1976) has long been known as a fundamental approach to multiple testing with familywise error rate (FWER) control. Recently, Xu et al. (2025) introduced an e-Closure Principle for FDR control. Given a local e-value $E_{S}$ for each intersection hypothesis $H_{0}^{S}=\bigcap_{i\in S}H_{0}^{i}$, $S\subseteq[p]$, meaning $\mathbb{E}_{H_{0}^{S}}[E_{S}]\leq 1$, we consider the closure set
@@
where $\boldsymbol{E}=(E_{S})_{S\subseteq[p]}$ and $\textnormal{FDP}_{S}(R)=\frac{|S\cap R|}{|R|\lor 1}$. Xu et al. (2025) showed that any rejection set $R\in\mathcal{R}_{\alpha}(\boldsymbol{E})$ controls the FDR and any procedure that controls the FDR at level $\alpha$ can be constructed in this way.
Furthermore, Xu et al. (2025) showed that these closure sets even permit a more sophisticated error rate control that goes beyond the classical notion of FDR control, i.e., $\alpha$ and $R\in\mathcal{R}_{\alpha}(\boldsymbol{E})$ can be chosen post-hoc! Given a family of e-values $\boldsymbol{E}=(E_{S})_{S\subseteq[p]}$, they proved that
$\mathbb{E}\left[\sup_{\alpha>0}\max_{R\in\mathcal{R}_{\alpha}(\boldsymbol{E})}\frac{\textnormal{FDP}_{I_{0}}(R)}{\alpha}\right]\leq 1.$ (3)
@@
Note that the post-hoc error rate control in (4) is slightly different from classical FDR control (where the FDR is bounded by $\alpha$), since $\tilde{\alpha}$ is inside the expectation. If $\tilde{\alpha}$ was fixed in advance, then the two notions would be equivalent due to the linearity of the expected value. However, since $\tilde{\alpha}$ is allowed to be a random variable depending on the data, the classical notion needs to be adapted by drawing $\tilde{\alpha}$ inside the expectation. This notion of post-hoc validity was also considered in (Grünwald, 2024; Koning, 2023) for single hypotheses.
@@
Example 1
Suppose all null hypotheses are true ($I_{0}=[p]$). In that case $\textnormal{FDP}_{I_{0}}(R)$ becomes $1$ if $R\neq\emptyset$ and 0 otherwise. Suppose we have a multiple testing procedure $R_{\alpha}$ for each $\alpha\in(0,1]$ that has exact error rate control, meaning $\mathbb{E}[\textnormal{FDP}_{I_{0}}(R_{\alpha})]=\alpha$ (for example, this is satisfied by the Benjamini-Hochberg procedure (Benjamini and Hochberg, 1995) with independent and uniformly distributed null p-values). Now set $\tilde{\alpha}=\inf\{\alpha\in(0,1]:R_{\alpha}\neq\emptyset]$. Then, we have